# Topology-Preserving Image Segmentation with Spatial-Aware Persistent Feature Matching

### v2

@@ 4.6 Ablation Studies on Weight $\lambda$
The results in Table 1, Table 2, Table 3 reports under the best weight $\lambda$ (Eq. 6) for the topological loss functions studied in this section. We experiment the WTLoss, He et al. and our proposed method using weights from 1e-5 to 1e-2. For TCLoss whose value is from fewer persistent features, we extend the range to 1e-0. For clDice and BMLoss, we use the optimal weights since they are reported in their papers (0.3 for clDice on CREMI, 0.5 on other datasets; 0.5 for BMLoss on both datasets). All experiments in the ablation study are also repeated with 3 random seeds and the mean value is reported in the charts. For our method and the two methods without spatial-aware matching (WTLoss and He et al.), as reported in Fig. 6, we observe that the topological accuracy is higher if a larger $\lambda$ is applied. However, if $\lambda$ is too large, the training will either converges badly or produce results with many small holes on the foreground that significantly increase the 1-Betti error (which we denote by dotted lines and crosses in Fig. 6). Our proposed method usually allows using a higher weight to further reduce the topological error. This effect is observed in repeated experiments over different datasets. A potential explanation is that the normalized spatial weights decrease the magnitude of gradient of the loss function. Nevertheless, even under the same weight, our method still consistently outperforms the others.
@@ 6.1 Roads
The (Massachusetts) Roads dataset [23] (Link) has in total 1171 $1500\times 1500$ images, and is split into 1108 training, 14 validation and 49 testing by the dataset author. We follow this dataset split in our experiments. Each image is cropped into 25 $300\times 300$ image patches. In total, there are 27700 for training, 350 for validation and 1225 for testing. The evaluation metrics are computed on the same patch size (the same for the following datasets).
6.2 CREMI
The CREMI dataset [9] (Link) has three volumes (A, B & C), volume A, B has 125 scans, volume C has 123 scans. Each scan is an $1250\times 1250$ 2D image. We use volume A,B and scan 1-14, 16-74, 76-81 in volume C for training, scan 82-92 in volume C for validation and scan 93-125 for testing. Each scan is cropped into 25 $250\times 250$ patches. In total, there are 8225 for training, 275 for validation and 825 for testing. The original CREMI dataset only provide the ground truth mask for each cell. We take the spaces between the cells to form the ground truth for the cell boundaries (extracellular matrix).
@@ 6.3 CrackTree
6.4 DRIVE
@@ 6.5 Roads-small
We follow as many details as we can in [30] to prepare the Roads-small dataset. In total 134 images are selected from the original Roads dataset (mixing training, validation and testing, since we use 3-fold cross validation for this dataset). In [30], which images are selected is not provided so we pick image by ourselves. It is difficult to describe the selected images unless providing a long list here but the general rule is to select images with more complexed road networks and less blank regions. The selected images are then divided randomly into training, validation and testing for each fold. The images are then downsampled from $1500\times 1500$ to $375\times 375$. Again, details of downsampling are not provided in [30], so we use (the same for C.Elegan-small) bicubic interpolation for all images and labels and then binarized the downsampled labels by thresholding them with a value of 0.5. No further manual correction is made to the downsampled labels. Before cropping, the images are padded to the right and bottom to $384\times 384$. Then, each image is cropped to 64 $48\times 48$ patches. In total, there are 6400 for training, 640 for validation and 1536 for testing.
6.6 C.Elegan-small
The C.Elegan dataset [31] (Link) has 200 $696\times 520$ images. We crop the images and only keep the center $340\times 340$ effective imaging area and then downsample them to $96\times 96$. For fold 1, we use A01-C22 for training, C23-D08 for validation and D09-E04 for testing; For fold 2, we use B07-E04 for training, A01-A10 for validation and A11-B06 for testing; For fold 3, we use A01-B16, C23-E04 for training, C13-C22 for validation and B17-C12 for testing. Each image is cropped into 4 $48\times 48$ patches. In total, there are 280 for training, 40 for validation and 80 for testing.
@@ 8 More Discussion on Experimental Results
Figure 7: More qualitative results to compare with BMLoss on Roads-small dataset. From left to right: image, ground truth, BMLoss, SATLoss.
@@ 9 Additional Results
Table 5 shows the full training time comparison. The ablation results on the other datasets are provided in Fig. 8, Fig. 9, Fig. 10, Fig. 11, Fig. 12, Fig. 13. Due to limited resources and for consistency between datasets, the ablation experiments were only run on the first fold. In addition, we provide more qualitative results in Fig. 14 for our main experiments. In Fig. 7, we also show more qualitative comparison with on the Roads-small dataset where our method fails to outperform the BMLoss in some metrics. In Fig. 15, we show more persistent feature matching results. In Fig. 16, we provide an illustrative example of the motivation of the spatial-aware matching, how it works during the matching process and how it helps with the topological accuracy of the segmentation.
@@
Roads 52k 2.9 44 43 44 $>$20000
CREMI 26k 1.0 16 16 16 $>$8000
CrackTree 3.7k 0.23 3.5 3.5 3.5 $>$1400
DRIVE 2.4k 0.03 0.22 0.23 0.23 $>$100

### v3

@@ 4.6 Ablation Studies on Weight $\lambda$
The results in Table 1, Table 2, Table 3 report under the best weight $\lambda$ (Eq. 6) for the topological loss functions studied in this section. We experiment the WTLoss, He et al. and our proposed method using weights from 1e-5 to 1e-2. For TCLoss whose value is from fewer persistent features, we extend the range to 1e-0. For clDice and BMLoss, we use the optimal weights since they are reported in their papers (0.3 for clDice on CREMI, 0.5 on other datasets; 0.5 for BMLoss on both datasets). All experiments in the ablation study are also repeated with 3 random seeds and the mean value is reported in the charts. For our method and the two methods without spatial-aware matching (WTLoss and He et al.), as reported in Fig. 6, we observe that the topological accuracy is higher if a larger $\lambda$ is applied. However, if $\lambda$ is too large, the training will either converges badly or produce results with many small holes on the foreground that significantly increase the 1-Betti error (which we denote by dotted lines and crosses in Fig. 6). Our proposed method usually allows using a higher weight to further reduce the topological error. This effect is observed in repeated experiments over different datasets. A potential explanation is that the normalized spatial weights decrease the magnitude of gradient of the loss function. Nevertheless, even under the same weight, our method still consistently outperforms the others.
@@ 6 Details on Datasets
6.1 CREMI
The CREMI dataset [9] (Link) has three volumes (A, B & C), volume A, B has 125 scans, volume C has 123 scans. Each scan is an $1250\times 1250$ 2D image. We use volume A,B and scan 1-14, 16-74, 76-81 in volume C for training, scan 82-92 in volume C for validation and scan 93-125 for testing. Each scan is cropped into 25 $250\times 250$ patches. In total, there are 8225 for training, 275 for validation and 825 for testing. The original CREMI dataset only provide the ground truth mask for each cell. We take the spaces between the cells to form the ground truth for the cell boundaries (extracellular matrix). The evaluation metrics are computed on the same patch size (the same for the following datasets).
6.2 DRIVE
@@ 6.3 Roads
The (Massachusetts) Roads dataset [24] (Link) has in total 1171 $1500\times 1500$ images, and is split into 1108 training, 14 validation and 49 testing by the dataset author. We follow this dataset split in our experiments. Each image is cropped into 25 $300\times 300$ image patches. In total, there are 27700 for training, 350 for validation and 1225 for testing.
@@ 6.4 CrackTree
6.5 C.Elegan-small
We follow as many details as we can in [31] to prepare the C.Elegan-small and Roads-small dataset. The C.Elegan dataset [32] (Link) has 200 $696\times 520$ images. We crop the images and only keep the center $340\times 340$ effective imaging area and then downsample them to $96\times 96$. For fold 1, we use A01-C22 for training, C23-D08 for validation and D09-E04 for testing; For fold 2, we use B07-E04 for training, A01-A10 for validation and A11-B06 for testing; For fold 3, we use A01-B16, C23-E04 for training, C13-C22 for validation and B17-C12 for testing. Each image is cropped into 4 $48\times 48$ patches. In total, there are 280 for training, 40 for validation and 80 for testing.
@@ 6.6 Roads-small
In total 134 images are selected from the original Roads dataset (mixing training, validation and testing, since we use 3-fold cross validation for this dataset). In [31], which images are selected is not provided so we pick image by ourselves. It is difficult to describe the selected images unless providing a long list here but the general rule is to select images with more complexed road networks and less blank regions. The selected images are then divided randomly into training, validation and testing for each fold. The images are then downsampled from $1500\times 1500$ to $375\times 375$. Again, details of downsampling are not provided in [31], so we use (the same for C.Elegan-small) bicubic interpolation for all images and labels and then binarized the downsampled labels by thresholding them with a value of 0.5. No further manual correction is made to the downsampled labels. Before cropping, the images are padded to the right and bottom to $384\times 384$. Then, each image is cropped to 64 $48\times 48$ patches. In total, there are 6400 for training, 640 for validation and 1536 for testing.
@@ 9 Additional Results
Table 5 shows the full training time comparison. The ablation results on the other datasets are provided in Fig. 7, Fig. 8, Fig. 9, Fig. 10, Fig. 11, Fig. 12. Due to limited resources and for consistency between datasets, the ablation experiments were only run on the first fold. In addition, we provide more qualitative results in Fig. 13 for our main experiments. In Fig. 14, we show more persistent feature matching results. In Fig. 15, we provide an illustrative example of the motivation of the spatial-aware matching, how it works during the matching process and how it helps with the topological accuracy of the segmentation.
@@
Roads 52k 2.9 44 43 44 >>20000
CREMI 26k 1.0 16 16 16 >>8000
CrackTree 3.7k 0.23 3.5 3.5 3.5 >>1400
DRIVE 2.4k 0.03 0.22 0.23 0.23 >>100